OpenStudy (anonymous):

square root of 40 over square root of 8

OpenStudy (anonymous):

What do you think you have to do.

OpenStudy (anonymous):

Are you trying to simplify it, orr?

OpenStudy (anonymous):

Multiply the 8 to rationalize the denominator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

Urmm. I mean. That's not really how I'd do it. O.o

OpenStudy (anonymous):

And it is kind of wrong. >.>

OpenStudy (zzr0ck3r):

\(\frac{\sqrt{40}}{\sqrt{8}}=\sqrt{\frac{40}{8}}=\sqrt{5}\)